OpenStudy (anonymous):

√27 + 2√5 - √75 can sumone show the steps to this please!!

OpenStudy (saifoo.khan):

\[-5\sqrt{5} + 3\sqrt{3}\]

OpenStudy (anonymous):

\[\sqrt{3\times9} + 2\sqrt{5} - \sqrt{3\times25}\] \[3\sqrt{3} + 2 \sqrt{5} - 5\sqrt{3}\] \[2 \sqrt{5} - 2 \sqrt{3}\]